Manuel Kauers Hagenberg, 19. März 2003 RISC-Linz

Exercises

1. What are the building blocks in the following sentence:

sprechen – speak morgen – tomorrow im = in + dem über – about Arbeit f – work

Franz spricht morgen im Forum über die Arbeit des Instituts.

For mathematicians: If a sentence has n building blocks, how many admissible word orders are conceivable?

2. In jokes and sayings, word order can be used for timing effects by putting the point at the very end of the sentence. In this sence, find the best possible translation of the following two sentences:^{1}

need – brauchen

chess computer – Schachcomputer, m already – schon know – kennen/wissen video recorder – Videorekorder, m machine – Maschine f smart – klug than – als human – Mensch, m

I don’t need a chess computer. I already know from my video recorder that machines are smarter than humans.

3. Connect the following connectives with their English counterpart.

aber dass obwohl oder und während weil wenn

although and because but if or that while

Which ones are missing?

4. There are two families of connectives. What is their difference? Illustrate the difference with an example sentence of your choice. For instance, use the connectives denn and weil.

1Taken from the german comedian Harald Schmidt; Hint: Put the video recorder at the end.